OpenStudy (anonymous):

Can you help me solve this? (x^3 + y^3) / (x - y) @campbell_st

OpenStudy (michele_laino):

please you have to perform the division between polynomials, which is also called "long division"

OpenStudy (anonymous):

Well I know.... but i get confused on how to do it properly...

OpenStudy (michele_laino):

please divide x^3 by x

OpenStudy (anonymous):

3?

OpenStudy (michele_laino):

what is your result?

OpenStudy (michele_laino):

sorry, division between power with the same base is another power with the same base and, for exponent, the difference between the exponent of dividend and divisor so x^3:x=x^(3-1)=... please complete

OpenStudy (anonymous):

I am not sure I understand what you are saying...

OpenStudy (anonymous):

x^2or-X^2 idk

OpenStudy (michele_laino):

ok! x^2, now multiply x^2 by the divisor, namely calculate this: \[x ^{2}(x+y)=...\] your first step is :